please help me solve the following problem.
2/3 + 1/4t = 1/3

Answer by stanbon(75887)

2/3 + 1/4t = 1/3
-----
Multiply thru by 12 to get:
8 + 3t = 4
3t = -4
t = -4/3
============
Cheers,
Stan H.